You bought six numbers at your local hardware store. The numbers are 0, 1, 2, 3, 4, 5.
a) How many 6 digit house numbers would be even?
(note: 0351 is not considered a house number and you only use each number once).
b) Using the same digits as above, how many 4 digit numbers smaller than 3500 could
you form if repetition was not allowed? (note: 0351 is not considered a house number).

Found 2 solutions by Edwin McCravy, Aldorozos:
Answer by Edwin McCravy(20054)

You bought six numbers at your local hardware store.
The numbers are 0, 1, 2, 3, 4, 5.
a) How many 6 digit house numbers would be even?
(note: 0351 is not considered a house number as you
only use each number once).

First we find the number of ways to choose the 1st
and 6th (last) digits.

Case 1.  We choose 2 or 4 for the first digit

We choose the first digit 2 ways, 2 or 4
We choose the last digit 2 ways, 0 and whichever 
of 2,4 we did not choose for the first digit.

That's 2*2 = 4 ways for case 1.

Case 2. We choose 1,3, or 5 for the 1st digit

We choose the 1st digit 3 ways, 1,3 or 5
We choose the 6th (last) digit 3 ways, 0, 2, or 4 
   choose for the first digit.

That's 3*3 = 9 ways for case 2.

Total number of ways to choose 1st and 6th (last) digits =
4+9 = 13 ways to choose 1st and 6th (last) digits.

For each of those 13 ways to choose the 1st and 6th digits,
we can choose the 2nd digit any of the 4 remaining ways.
we can choose the 3rd digit any of the 3 remaining ways.
we can choose the 4th digit either of the 2 remaining ways.
we can choose the 5th digit only 1 remaining way.

Answer:  13*4*3*2*1 = 312 ways.

----------------------------

b) Using the same digits as above, how many 4 digit
numbers smaller than 3500 could you form if repetition
was not allowed? (note: 0351 is not considered a house
number).
This does not say anything about the number being even or odd,
so we will assume it can be either even or odd.

First we find the number of ways to choose the 1st
and 2nd digits.

Case 1. The first digit is 3
        We choose the first digit as 3.  That's 1 way.
        We choose the 2nd digit as 0, 1, 2, or 4.  That's 4 ways.
That's 1*4 = 4 ways for case 1.

Case 2. The first digit is 1 or 2
        We choose the first digit as 1 or 2.  That's 2 ways
        We choose the 2nd digit as any of the 5 remaining digits.
That's 2*5 = 10 ways for case 2.

Total for cases 1 and 2 = 4+10 = 14

For each of those 14 ways to choose the 1st and 2nd digits,
we can choose the 3rd digit any of the 4 remaining ways.
we can choose the 4th digit any of the 3 remaining ways.

Answer: 14*4*3 = 168 ways

Edwin

Answer by Aldorozos(172)

Here is to the best of my knowledge
Here is the house number
NNNNNN (these are six numbers)
I am assuming that by saying 0351 is not considered a house number means a house number can't start with number zero.
The house number can be even if it ends with 0, 2 , 0r 4
This means it should look like one of the three forms below
NNNNN0
NNNNN2
NNNNN4

No we look at NNNNN0
If that is the came for the first N in our house number NNNNN we have 5 options to choose from. These options are 1,2,3,4,5. The last one is supposed to be zero to get an even number. This means we can choose any of the six numbers except 0 for the first N. Which means we can choose any of the five numbers.
For the second N we have four options since we already used zero as the last number and we used one of our numbers as the first number. For the third N we have three options, for the fourth we have 2 options for the fifth we have 1 option.
Therefore our options for NNNNN0 = 5*4*3*2*1 = 120 options
For NNNNN2 we have 4*4*3*2*1 = 98 (the first N can be any number except 2 and zero. since 2 is used at the end to make the number even and zero can't be used)
Similarly for NNNNN4 we have 4*4*3*2*1 = 98
Now we add the three numbers = 120+98+98 = 316


Part B for numbers to be less than 3500. The house number is like NNNN
The highest number we can have is 3452. This means that the first number has to be 3 or less and it can't be 0. This means the first number can be 1, or 2 or 3. Therefore we have three options for the first number. The reason we are saying in can't be zero is because the problem says the house number can't be 0351 which we assume that it means it can't start with zero.
For the second number in 3453, we can have 4 or any number less than 4 which means 0, 1,2,3 or 4. However we have used 1, or 2, or 3 as the first digit. Therefore we only have four numbers left. This means we have four options for the second number.
For the third this digits in 3453, we have number 5. We can have six digits for this option 0,1,2,3,4,5. However we have used up two of the digits for the firs two digits. Therefore we are left with only 4 letters (options)
For the last digits in 3452 (number 2) we can have three options since we have used up three of the numbers. Please note that we have a number like 3425. This number is less than 3500 and is also less than 3452
Now we multiply out options for the first N and second N, third N and the fourth N 3*4*4*3 = 144